stats question

A study conducted at a certain college shows that 55% of the school's graduates find a job in their chosen field within a year after graduation. Find the probabilitythat among 7 randomly selected graduates, at least one finds a job in his or her chosen field within a year of graduating.
Answer #1
n = 7 P(find a job) = 0.55, q = 1- P = 0.45 P(at least one) = 1- P(none) = 1- (0.45)^7 = 1 - 0.0037 = 0.9963
answered by: dreyvin

> what is the formula name? the 1-p^n

Emmanuel Perez Wed, Feb 23, 2022 12:19 PM
Answer #2

at least one gets a job = 1-nobody gets a job

=1-(45/100)^7
=1 - 0.00373669453 = 0.996263305
Answer #3
P(find a job) = .55
P(don't find a job) = 1-P(find a job) = .45
P(none find a job) = P(don't find a job) ^ 7 = .0037
P(at least one finds a job) = 1-P(non find a job) = .9963
answered by: I am not joe